A generator consists of a rectangular coil 84 cm by 1.4 m , spinning in a 0.15 T magnetic field.

If it's to produce a 60-Hz alternating emf with peak value 6.5 kV , how many turns must it have?
Answer #1

Peak emf = NBAW

Angular frequency = 2 × pi x f = 2× 3.14 × 60= 376.8

6.5× 10^3= N × 0.15 × 0.84 × 1.4 × 376.8

N= 98turns